OpenStudy (this):

some one pllzzz help me factor and each of the following quadratic equations 16b^2-25=0

OpenStudy (anonymous):

transpose - 25 to Right Hand Side 16b^2 = 25 then divide both sides by 16 b^2 = 25/16 can u solve now???

OpenStudy (anonymous):

@this Did you understand what I replied

OpenStudy (this):

no really but i will try

OpenStudy (anonymous):

have u understood the two steps that i have given??

OpenStudy (anonymous):

transpose - 25 to Right Hand Side 16b^2 = 25 then divide both sides by 16 b^2 = 25/16 now take square root of both sides, we get b = +/- sqrt(25)/sqrt(16) = +/- 5/4 or (+/- 1.25 in decimal form)

hartnn (hartnn):

16b^2 can be written as (...)^2 ? what comes inside brackets?

OpenStudy (anonymous):

16b^2-25=0 add 25 on both side(or the left hand side and the right hand side) 16b^2-25+25=0+25 16b^2=25 divide 16 0n both sides we get 16/16b^2=25/16 b^2=25/16 now square root both sides to get (b^2)^1/2=(25/16)^1/2 or \[\sqrt{b ^{2}}=\sqrt{\frac{ 25 }{ 16 }}\] \[b=\frac{ \sqrt{25} }{ \sqrt{16} }\] \[b=\frac{ 5 }{ 4 }\]

hartnn (hartnn):

why is ever1 solving it without factoring, when the question clearly says to factor it :\

OpenStudy (anonymous):

now that we know b=5/4 (b-5/4)(b+5/4)=0 multiply both sides by 4 (4b-5)(4b+5)=0

OpenStudy (anonymous):

from b =+-sqrt(25/16) b=+- 5/4 now that we know b=-5/4 and +5/4 (b-5/4)(b+5/4)=0 multiply both sides by 4 (4b-5)(4b+5)=0